The Alvin Secretarial Service procures temporary office personnel for major corporations. They have found that 60% of their invoices are paid within ten working days. A random sample of 19 invoices is checked. What is the probability that exactly 5 of the invoices will be paid within ten working days? Round your answer to four decimal places.
Answer #1

Here, n = 19, p = 0.6, (1 - p) = 0.4 and x = 5
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 5)
P(X = 5) = 19C5 * 0.6^5 * 0.4^14
P(X = 5) = 0.0024